# Green Edge AI: A Contemporary Survey

### v1

@@ VI-B Hardware/software Co-design
The existing methods for energy-efficient edge AI, including those reviewed in previous sections, mainly focus on software designs, and their benefits are greatly affected by the underlying hardware implementations [8]. It is hence critical to consider the hardware/software co-design for edge AI. A precursory study in [259] suggests to deploy multiple pruned CNNs instead of an unpruned one for inference, which not only increases the robustness and accuracy, but also offers the opportunities to use ultra-low-power processing cores and scale-down the memory access voltage. Inspired by NAS, hardware-aware NAS (HW-NAS) was proposed by factoring the hardware-related metrics (e.g., latency and energy consumption) and constraints (e.g., bit-width and memory) in the search objectives [260]. However, it confines the search space to a pool of candidate neural network architectures, and the search results are only suitable for fixed hardware configurations. Given the uprise of reconfigarable AI accelerators such as the field programmable gate arrays (FPGAs), expanding the search space of HW-aware NAS by incorporating different configurations of hardware platforms is a natural extension. For example, [261] models the hardware search space as an FPGA pool and develops a two-level exploration frameowork. The experimental result shows that with hardware/software co-design, the EE achieves a remarkable $54.05\%$ increase on ImageNet classification with Xilinx FPGAs. In addition, cross-layer optimization from hardware, software, to complier, can be conducted to further improve EE [262]. The significant challenges of hardware/software co-design include the huge search space and cost of searching algorithms, given diverse edge AI hardware platforms [263, 264]. It is also critical for hardware and software researchers of edge AI to collaborate and establish clear understanding on design objectives and constraints outside their own expertise [265].
@@ VI-C Neuromorphic Computing
Human brains are among the most energy-efficient computing devices, which can provide $10^{18}$ FLOPS with merely $20$ Watts of power consumption [266]. Neuromorphic computing endeavors to imitate operations in human brains, where biological neurons communicate with electrochemical signals through action potentials [267]. SNNs are a family of neuromorphic algorithms. They differs from the conventional DNNs by encoding input as time-series and use discrete spikes to transfer information between spiking neurons, each of which can fire a spike with sufficient input strength called membrane potential [268]. Thus, SNNs are orders of magnitude more energy-efficient than DNNs especially with the support of neuromorphic chips, such as the IBM’s TrueNorth and Intel’s Loihi chips. As a proof-of-concept, Accenture Labs demonstrated via an automotive voice control application that, an SNN deployed at Intel’s Kapoho Bay, which is a computing stick empowered by Intel’s Loihi chips, outperforms a CNN running at a GPU by 1000×1000$\times$1000 × and $200$ ms in terms of EE and latency, respectively, while maintaining comparable inference accuracy [269]. SNNs, however, bring new challenges in training [270] and hardware implementation [271]. Notably, there is also recent interest of training SNNs under the framework of FEEL [272, 273].
Neuromorphic computing also deviates from the traditional von Neumann architectures with separate units for processing and memory, where the energy spent on fetching data from off-chip memory for DNN computations is significant. Hence, it inspires to adopt the compute-in-memory (CIM) techniques, which perform MAC operations in the analog domain within memory subarray [274]. The history of CIM starts from the 1990s, and it receives great attention since the last decade with the developments of transistor and non-volatile memory technologies [275]. In 2021, an analog matrix processor for high-performance AI inference based on static random access memory (SRAM) was announced by ${}^{\circledR}$, which delivers 25 tera operations per second (TOPS) but just needs 3 Watts of power, i.e., only $1/10$ of the desktop GPU power consumption [276].
@@ VI-D Green Energy-powered Edge AI
Green energy sources such as solar, wind, and ambient radio frequency (RF) energy that can be scavenged by energy harvesting technologies, are key drivers for energy-efficient IoT applications [277]. Performing DNN inference on green energy-powered EUDs thus becomes a perfect candidate to achieve energy-efficient edge inference. However, given the intensive computation workloads, the intermittent green energy availability may frequently interrupt the inference processes, which need to be restarted when sufficient energy is collected [278]. To tackle this fundamental challenge, the task-based execution model from intermittent computing was adapted for DNN inference in [279] to build a full-scale RF energy-powered prototype. Since the available green energy is time-varying, adaptive inference is essential to maximize the long-term average accuracy. This idea was grounded by deploying a multi-exit DNN in [280], where the best exit is selected via Q-learning in runtime according the available energy level. Nevertheless, these works concentrate on on-device inference, and methods to efficiently exploit computational resources at the MEC server remains uncharted.
@@
Compared with green energy-powered edge inference, green energy-powered edge learning is much less investigated. Recently, FEEL systems where the EUDs are supported by green energy sources were proposed [281, 282]. Likewise, due to the unstable energy availability, EUDs might drop out of the training processing unexpectedly, rendering model convergence hard to guarantee with conventional FEEL algorithms. Fortunately, it was showed in [281] that such an issue can be overcome by re-scaling the learning rate with the energy arrival probability when the training loss function is strongly convex. Besides, green energy management is also a crucial aspect that was investigated in [282], where the number of participating green energy-powered EUDs was maximized via joint user scheduling and power allocation. However, a comprehensive understanding on the general convergence behavior of green energy-powered FEEL is still missing, and learning- and green energy-aware resource management strategies are pending developments.
@@ VII Conclusions
In an era with the proliferation of edge devices and surging demand for real-time intelligent processing, the energy consumption associated with edge AI applications poses not only grand operational challenges but also increasing environmental concerns. Understanding and mitigating the energy footprint of the related technologies thus become imperative for building energy-efficient and eco-friendly edge AI systems.
This paper delves into the critical realm of green edge AI, concentrating on the intersection of energy efficiency and AI at the edge. Our exploration has provided a comprehensive overview of strategies to mitigate the formidable energy challenge of the emerging edge AI landscape. From dissecting the energy consumption in basic tasks of edge AI such as data acquisition, edge learning, and edge inference, to unraveling the design principles that emphasize energy efficiency, adaptability, and the judicious trade-off between intelligence and greenness, this paper navigates through the multifaceted dimensions of green edge AI. Promising future research directions, including integrated sensing and communication, hardware/software co-design, neuromorphic computing, and green energy-powered edge AI, have also been highlighted to attract early attention. In conclusion, it is evident that the synthesis of technological advancements, algorithmic innovations, and strategic energy management is crucial toward realizing the 6G vision of sustainable and pervasive intelligence. We anticipate that this survey paper will serve as a valuable reference for researchers, practitioners, and policymakers engaged in shaping the trajectory of green edge AI.

### v2

@@ VI-B Hardware/software Co-design
The existing methods for energy-efficient edge AI, including those reviewed in previous sections, mainly focus on software designs, and their benefits are greatly affected by the underlying hardware implementations [8]. For example, implementing DNN inference applications on field programmable gate arrays (FPGAs) and application specific integrated circuits (ASICs) are more energy-efficient than on general-purpose CPUs and GPUs [285]. Also, a recent study in [286] found that even for smartphones of the same brand, the energy consumption and AI performance can be very different for certain DNN inference tasks. It is hence critical to investigate the hardware/software interplay for improving EE of edge AI.
A precursory study in [287] suggests to deploy multiple pruned CNNs instead of an unpruned one for on-device inference, which not only increases the robustness and accuracy, but also offers the opportunities to use ultra-low-power processing cores and scale-down the memory access voltage for energy savings. Inspired by NAS, hardware-aware NAS (HW-NAS) was proposed by factoring the hardware-related metrics (e.g., latency, energy consumption, and memory space) in the search objectives and constraints [288]. However, it normally confines the search space to a pool of candidate neural network architectures, and the search results are only suitable for fixed hardware configurations. Given the popularity of reconfigurable AI accelerators such as FPGAs, expanding the search space of HW-aware NAS by incorporating different configurations of hardware platforms is a natural extension. For example, [289] models the hardware design space as an FPGA pool and develops a two-level exploration framework. The experimental result shows that with hardware/software co-design, the EE achieves a remarkable $54.05\%$ increase on ImageNet classification with Xilinx FPGAs. In addition, cross-layer optimization from hardware, software, to compiler, can be conducted to further improve EE [290]. The significant challenges of hardware/software co-design include the huge search space and cost of searching algorithms, given diverse DNN architectures and edge AI hardware platforms [291, 292]. To resolve these challenges, it is critical for hardware and software researchers of edge AI to collaborate and establish clear understanding on design objectives and constraints outside their own expertise as suggested in [291, 293]. Moreover, as DNN training has different computation and memory access patterns compared to DNN inference [294, 295], training-specialized hardware/software co-design strategies are needed to maximize the EE of cooperative edge learning.
@@ VI-C Neuromorphic Computing
Human brains are among the most energy-efficient computing devices, which can provide $10^{18}$ floating point operations per second with merely $20$ Watts of power consumption [296]. Neuromorphic computing endeavors to imitate operations in human brains, where biological neurons communicate with electrochemical signals through action potentials [297]. SNNs are a family of neuromorphic algorithms. They differ from the conventional DNNs by encoding input as time-series and using discrete spikes to transfer information between spiking neurons, each of which can fire a spike with sufficient input strength called membrane potential [298]. Thus, SNNs are orders of magnitude more energy-efficient than DNNs especially with the support of neuromorphic chips, such as the IBM’s TrueNorth and Intel’s Loihi chips. As a proof-of-concept, Accenture Labs demonstrated via an automotive voice control application that, an SNN deployed at Intel’s Kapoho Bay, which is a computing stick empowered by Intel’s Loihi chips, outperforms a CNN running at a GPU by 1000×1000$\times$1000 × and $200$ ms in terms of EE and latency, respectively, while maintaining comparable inference accuracy [299]. Notably, there are recent interests of training SNNs under the framework of FEEL [300, 301] and SNN-based cooperative edge inference [302]. However, SNNs bring new challenges in training [303] and hardware implementation [304]. Latest studies have also found that EE improvements achieved by SNNs over DNNs vary with inference task types [305, 306]. Hence, properly selecting between SNNs and DNNs based on application needs is crucial.
Neuromorphic computing also deviates from the traditional von Neumann architectures with separate units for processing and memory, with which, the energy spent on fetching data from off-chip memory for DNN computations is significant. Hence, it inspires to adopt the CIM techniques for edge AI, which can perform MAC operations in the analog domain within memory subarray [307]. The history of CIM starts from the 1990s, and it receives great attention since the last decade with the developments of transistor and memory technologies [308]. In 2021, an analog matrix processor for high-performance AI inference based on non-volatile flash memory was announced by Mythic®, which delivers 25 tera operations per second (TOPS) but just needs 3 Watts of power, i.e., only $1/10$ of the desktop GPU power consumption [309]. However, limitations of current CIM technologies, e.g., sensitive to analog errors [310] and immature development software [311], await to be addressed.
@@ VI-D Green Energy-powered Edge AI
Although methods to reduce the non-renewable energy consumption are paramount in no doubt, green energy sources such as solar, wind, and ambient radio frequency (RF) energy that can be scavenged by energy harvesting technologies, are new drivers for energy-efficient IoT applications [312]. Performing DNN inference on green energy-powered EUDs thus becomes a perfect candidate to achieve energy-efficient edge inference. However, given the intensive computation workloads, the intermittent green energy availability may frequently interrupt the inference processes, which need to be restarted when sufficient energy is collected [313]. To tackle this fundamental obstacle, the task-based execution model from the area of intermittent computing was adapted for DNN inference in [314] to build a full-scale RF energy-powered prototype. Since the available green energy is time-varying, adaptive inference is essential to maximize the long-term average accuracy. This idea was first grounded by deploying a multi-exit DNN in [315], where the best exit is selected via Q-learning in runtime according to the level of available energy. Green energy-aware adaptive inference was also achieved by selecting models with different energy overhead and accuracy performance [316]. Nevertheless, these works concentrate on on-device inference, and methods to efficiently exploit the computational resources at the MEC server via device-edge co-inference remains uncharted.
@@
Compared with green energy-powered edge inference, green energy-powered edge learning is much less investigated. Recently, FEEL systems where the EUDs are supported by green energy sources were proposed in [317, 318]. Likewise, due to the unstable energy availability, EUDs might drop out of the training process unexpectedly, rendering model convergence hard to guarantee with conventional FEEL algorithms such as FedAvg. Fortunately, it was showed in [317] that such an issue can be overcome by re-scaling the learning rate with the green energy arrival probability when the training loss function is strongly convex. Besides, green energy management is also a crucial aspect that was investigated in [318], where the number of participating green energy-powered EUDs was maximized via joint user scheduling and power allocation. However, a comprehensive understanding on the general convergence behavior of green energy-powered FEEL is still absent. Also, joint learning- and green energy-aware resource management strategies are pending development. Beyond vanilla FEEL, it is interesting to integrate green energy sources with other emerging distributed edge learning paradigms, such as DEEL and H-FEEL discussed in Section IV-D.
VI-E Green Generative Edge AI
GenAI techniques, which can create contents such as text, images, speech, and videos [319], are unleashing the next wave of AI innovations, especially after the release of ChatGPT by OpenAI in late 2022. At the moment, many impressive GenAI applications are underpinned by foundation models (FMs) that have a few to hundreds of billions of parameters and need to be trained on massive datasets [320]. Apparently, training such giant models consumes huge amount of computational and energy resources, and thus it can merely take place at large datacenters. Although with less resource demand, making inference via FMs for GenAI services remains mostly in the Cloud, with prompts being uploaded by EUDs. Similar to other intelligent mobile applications, migrating GenAI services to mobile edge networks, namely generative edge AI, can benefit EUDs with low latency and immersive experience [321]. However, by the large-size nature of FMs, it is critical to fit the compute-intensive GenAI workloads to the resource-limited MEC environments and improve the EE.
Although it is unlikely to train FMs from scratch within mobile edge networks, fine-tuning pre-trained models with the data and computational resources at EUDs is plausible for personalization and task adaptation. FEEL is promising for this purpose [322]. However, updating billions of parameters at once may encounter the memory wall, e.g., fine-tuning the LLaMA model [323] even with a small batch size of 8 incurs more than 20 GB of peak memory usage [324], not to mention the significant on-device energy consumption and processing latency. Hence, parameter-efficient fune-tuning methods are necessarily, which freeze most model parameters to minimize the on-device computation and communication overhead. To further reduce the on-device computation energy consumption or with extremely restrictive on-device computational resources, the split learning (SL) framework can be applied [325]. For FM-based inference, existing model compression methods can be adapted for energy-efficient deployments at EUDs or MEC servers [326], which should preserve the versatility and generalizability of FMs. Another approach to reduce the inference energy cost is to promote device cooperation motivated by the fact that requested contents from nearby EUDs are usually correlated. This approach was demonstrated via diffusion model-based image generation in [327], where the shared processing steps are computed on the MEC server while EUDs just need to handle the prompt-specific computations. Moreover, energy-efficient SNNs have recently been found capable of language generation and understanding in [328] for edge applications. Although still in the conceptual stage, green generative edge AI is surely an area worthy of attention given the intense conflict between significant computation overhead of GenAI workloads and limited energy availability at EUDs.
@@ VII Conclusions
In an era with the proliferation of edge devices and surging demand for real-time intelligent processing, the significant energy consumption associated with edge AI applications poses not only grand operational challenges but also steadily increasing environmental concerns. Understanding and mitigating the energy footprint of the related technologies thus become imperative for building energy-efficient and eco-friendly edge AI systems. This paper delves into the critical realm of green edge AI, concentrating on the intersection of energy efficiency, AI, and edge computing. Our exploration has provided a comprehensive overview of strategies to mitigate the formidable energy challenge of the emerging edge AI landscape. From dissecting the energy consumption in basic tasks of edge AI such as data acquisition, edge learning, and edge inference, to unraveling the design principles that emphasize energy efficiency, adaptability, and the judicious trade-off between intelligence and greenness, this paper navigates through the multifaceted dimensions of green edge AI. Promising future research directions, including integrated sensing and communication, hardware/software co-design, neuromorphic computing, green energy-powered edge AI, and green generative edge AI, have also been highlighted to attract early attention from both academia and industry. In conclusion, it is evident that the synthesis of technological advancements, algorithmic innovations, and strategic energy management is crucial toward realizing the 6G vision of sustainable and pervasive intelligence. We anticipate that this survey paper will function as a timely and valuable reference for researchers, practitioners, and policymakers engaged in shaping the trajectory of green edge AI.